# Luminance-Aware Statistical Quantization: Unsupervised Hierarchical Learning for Illumination Enhancement

### v1

@@ 1 Introduction
In low-light environments, images frequently experience degradations like reduced visibility and heightened noise, which hinder subsequent vision-related tasks [1, 2, 3, 70]. Low-light image enhancement (LLIE) aims to reconstruct perceptually natural scenes by establishing mappings between low-light and normal-light distributions. However, this problem is fundamentally ill-posed, as natural luminance transitions follow continuous physical processes governed by scene radiance and sensor responses, rather than discrete pixel-level correspondences.
While recent deep learning methods—whether supervised [10, 32, 66, 65, 67] or unsupervised [11, 69, 68]—attempt to model light variations through paired or unpaired training, they inherently overfit to static relationships between low/normal-light domains. Supervised methods rely on pixel-level correspondences in paired data, forcing models to prioritize localized correlations over the physics of gradual luminance evolution. Unpaired approaches, though avoiding direct pairing, still depend on pseudo-references derived from empirical gamma corrections [11], inheriting prior biases. Both paradigms oversimplify the inherently context-dependent and continuous nature of luminance dynamics, resulting in constrained generalization: models excel in domain-specific scenarios with reference normal-light samples but struggle to adapt to unseen environments or sensor-specific degradations [4, 5]. This underscores the necessity for a paradigm shift toward learning luminance transitions from the intrinsic continuity of real-world illumination processes.
@@ 2.1 LLIE via Pixel-Level Consistency
Contemporary approaches focusing on pixel-level consistency can be categorized into three evolutionary stages [10, 8, 62, 9, 60, 58, 63, 56]. Early works like LLNet [34] and Retinex-Net [36] leveraged paired datasets to train CNNs with pixel-wise losses, achieving precise local corrections but suffering from domain overfitting. Methods like EnlightenGAN [42] introduced cycle-consistency constraints, while Zero-DCE [39] used non-paired training with empirical illumination curves, both inheriting biases from heuristic priors [2]. Recent diffusion models [11] enhanced flexibility through noise-to-clean transitions, and FeatEnHancer [1] proposed hierarchical feature fusion to bridge pixel-level and semantic gaps. These methods improved generalization but remained black-box mappings [3]. While progressively reducing dependency on strict pixel correspondences, these methods universally prioritize pixel-wise fidelity over the continuous, context-dependent nature of luminance transitions, resulting in constrained generalization when facing unseen scenarios or sensor-specific degradations [4].
@@ 2.2 LLIE with Illumination Priors
To alleviate these issues, several studies [39, 44, 7] have incorporated illumination-aware correction into deep neural networks, framing LLIE as a curve-estimation problem through different gamma correction samples. Specifically, [7] introduced learnable gamma transforms for illumination adjustment, yet enforced uniform corrections across all pixels. Methods like KinD [44] decomposed images into illumination-reflectance components, but relied on manually designed reflectance priors that oversimplified real-world radiance interactions, and its reliance on downstream pre-training objectives (e.g., perceptual losses) further introduces external biases that impair adaptability to diverse degradation patterns. More latest works like LightenDiffusion [11] integrated Retinex theory into diffusion steps, while GPP-LLIE [64] embedded illumination gradients as diffusion guidance. Though enhancing physical plausibility, these methods still imposed global correction strategies through rigid equation constraints.
@@ 3.3 Hierarchically‐Guided Diffusion
Forward Process with Hierarchical Guidance: The sampled set $\mathcal{H}=\{\mathcal{I}_{H}^{(n)}\}_{n=1}^{N}$ employs stochastic learning via diffusion transitions, exploiting the Markov property—each $\mathcal{I}_{H}^{(n)}$ relies only on its forerunner—to adaptively direct noise injection. The low-light image $\mathcal{I}_{L}$ and $\mathcal{H}$ are encoded together using $\mathcal{E}(\cdot)$, incorporating $k$ residual blocks and max-pooling for latent features $\mathcal{F}_{L}\in\mathbb{R}^{\frac{H}{2^{k}}\times\frac{W}{2^{k}}\times C}$ and $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}\in\mathbb{R}^{N\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}\times C}$. We align the $T$-step diffusion with $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}$ using a temporal mapping $\psi:\{1,\ldots,T\}\to\{1,\ldots,N\}$, $N\leq T$, by $\psi(t)=\lfloor t\cdot N/T\rfloor$. The forward diffusion process is reformulated as:
$q(x_{t}|x_{t-1},\mathcal{F}_{H}^{(\psi(t))})=\mathcal{N}\left(\sqrt{1-\beta_{t}}x_{t-1}+\tau_{t}\Delta\mu_{t},\beta_{t}I\right),\quad\Delta\mu_{t}=\mathcal{F}_{H}^{(\psi(t))}-x_{t-1},$ (7)
where $\Delta\mu_{t}$ represents the semantic shift induced by hierarchical guidance, and $\tau_{t}$ is a learnable adaptive weight. By analytically aggregating noise injections, the marginal distribution of $x_{t}$ becomes:
$q(x_{t}|x_{0},\mathcal{F}_{H}^{(\psi(t))})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sum_{s=1}^{t}w_{t,s}\left(\mathcal{F}_{H}^{(\psi(s))}-x_{0}\right),(1-\bar{\alpha}_{t})I\right),$ (8)
where $x_{0}=\mathcal{F}_{H}^{(n)}$ and $w_{t,s}=\sqrt{\bar{\alpha}_{t}}\frac{\tau_{s}\sqrt{1-\bar{\alpha}_{s-1}}}{\sqrt{\bar{\alpha}_{s}}}$ is determined by diffusion schedulers. This closed-form solution indicates that hierarchical labels progressively refine the diffusion trajectory through weighted summation, enabling coarse-to-fine semantic injection.
@@
where the $\mathcal{{D}}(\cdot)$ denotes the decoder. During inference, under the guidance of the low-light input $\mathcal{F}_{L}$, we employ the diffusion model’s implicit sampling strategy [18] for reverse denoising. The model utilizes its learned distribution to fit the optimal illumination-enhanced feature representation $\hat{\mathcal{F}}_{N}$, which is then decoded to yield the final output $\hat{\mathcal{I}}_{N}=\mathcal{D}(\hat{\mathcal{F}}_{N})$.
@@ 4.1 Experimental Settings
All experiments are carried out on a cluster of four NVIDIA A800 GPUs under Python 3.9 and PyTorch 2.0, with a fixed batch size of $16$. We employ the Adam optimizer [48], setting the denoising diffusion process learning rate to $2\times 10^{-5}$, while using a sampling ratio $k=3$. The hyperparameters $\lambda_{\text{d}}$, $\lambda_{\text{g}}$ and $\lambda_{\mathrm{GAN}}$ (if activated) are set to $0.9$, $0.005$ and $0.7$ respectively. Noise estimation during diffusion training is performed using the U-Net [49] architecture with $T=1000$ time steps.
@@ 4.2 Datasets and Metrics
We validated our approach using both paired and unpaired low-light benchmarks. For the paired evaluation, we used the LOLv1 [34] and LSRW [50] test sets—each comprising matched low- and normal-illumination image pairs—and reported restoration fidelity via PSNR and SSIM [52], alongside the full-reference perceptual score LPIPS [53]. To assess performance in the absence of ground-truth references, we then tested on the unpaired LIME [29], DICM [23], NPE [30], and VV [51] collections, measuring perceptual quality with the no-reference NIQE [54] and PI [55] metrics. Our comparative study encompasses six supervised approaches (RetinexNet [34], KinD++ [36], LCDPNet [56], URetinexNet [35], SMG [9] and PyDiff [57]) and six unsupervised approaches (Zero-DCE [39], EnlightenGAN [42], SCI [58], PairLIE [59], SCL-LLE [61], LightenDiffusion [11] and NeRCo [60].
@@ 4.3 Qualitative Comparison
As shown in Fig. 3, LASQ achieves enhanced local brightness adaptation and superior detail fidelity comparable to supervised methods URetinexNet [35] and KinD++ [36] on ground truth-annotated datasets, while demonstrating improved domain-adaptive color reproduction through integration with unpaired normal-light images in LASQ++. By contrast, existing methods exhibit distinct limitations: SCI [58] and SCL-LLE [61] suffer from persistent underexposure, whereas EnlightenGAN [42] produces blurred structural details, and NeRCo [60] tends to generate localized over-exposure artifacts. Furthermore, Fig. 4 confirms LASQ’s exceptional performance in real-world scenarios through its complete avoidance of local overexposure, noise amplification, and artifacts that persistently affect other methods: EnlightenGAN [42], NeRCo [60] and PairLIE [59] notably show severe localized overexposure and lens flare artifacts, and even supervised approaches URetinexNet [35] and KinD++ [36] still struggle to fully suppress these issues. Crucially, LASQ maintains natural scene characteristics without compromising detail fidelity or color consistency, thereby demonstrating unprecedented cross-scenario generalization capability across both constrained laboratory settings and unconstrained environmental conditions. This comprehensive evaluation systematically validates LASQ’s technical superiority in terms of adaptive illumination control, artifact suppression, and domain transfer effectiveness. More results will be provided in the Appendix.
@@ 4.4 Quantitative Comparison
The quantitative evaluation results across diverse datasets are summarized in Table 1, where LASQ demonstrates performance parity with leading supervised techniques on LOLv1 [34] and LSRW [50] while achieving state-of-the-art results among unsupervised methods through integration of unpaired normal-light images. Notably, on datasets DICM [23], NPE [30], and VV [51], LASQ outperforms existing approaches across most perceptual metrics, thereby confirming its intrinsic generalization prowess without domain-specific adaptation. Although normal-light reference integration improves color fidelity, its tendency toward overfitting partially counteracts the model’s inherent generalization capacity, resulting in metric degradation in LASQ++. Crucially, this performance highlights LASQ’s fundamental advantage in balancing domain adaptation with cross-scenario robustness, whereas LASQ++ prioritizes target-domain color accuracy at the expense of slight generalization capability. The extended experimental results, including comprehensive qualitative analyses and quantitative evaluations, are provided in the Appendix.
@@ 5.1 Fixed Luminance Adjustment
We replace our adaptive MCMC-based luminance adaptation with static, hand‐crafted functions (e.g., global gamma correction or fixed tone curves) drawn from prior LLIE methods [20, 25, 26, 27]. Embedding these into the forward diffusion degrades PSNR, SSIM, and LPIPS (Table 3), showing reduced feature richness and generative fidelity (Fig. 5). In contrast, our physics‐driven, multi‐scale operator—via power‐law stratification and adaptive sampling—better balances global consistency and local detail, yielding stronger cross‐scenario generalization and perceptual quality.
@@ Appendix A The Detailed LASQ Workflow
Hierarchically-Guided Diffusion: The forward process employs a hierarchically-guided diffusion framework to progressively enhance low-light images through semantically-aligned noise injection. By encoding the input $\mathcal{I}_{L}$ and its luminance-enhanced variants $\mathcal{H}$ into latent features $\mathcal{F}_{L}$ and $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}$, the method establishes temporal correspondence between diffusion steps and multi-scale guidance via a mapping function $\psi(t)$. This alignment dynamically steers the noise trajectory, where each transition step incorporates hierarchical semantic shifts $\Delta\mu_{t}$ derived from $\mathcal{F}_{H}^{(\psi(t))}$ to modulate the degradation path. Throughout the forward pass, coarse-to-fine illumination characteristics are gradually imprinted into the diffusion process through weighted aggregation of hierarchical guidance signals. During reverse denoising, the network $\epsilon_{\theta}$ iteratively removes noise while synergizing hierarchical cues from $\mathcal{F}_{H}^{(\psi(t))}$ and original low-light features $\mathcal{F}_{L}$. The denoising trajectory is regularized by a dual objective: a noise prediction loss ensures faithful reconstruction aligned with hierarchical semantics, while a global constraint $\mathcal{L}_{\text{g}}$ enforces the structural consistency.
@@ Appendix B The Algorithmic Framework
13: Define $\psi(t)=\lceil tN/T\rceil$, set $x_{0}=\mathcal{F}_{H}^{(1)}$
@@
16:  Get hierarchy feature $\mathcal{F}_{H}^{(\psi(t))}$
17:  $x_{t}\sim\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1}+\tau_{t}\Delta\mu_{t},\beta_{t}I)$

### v2

@@ 1 Introduction
In low-light environments, images frequently experience degradations like reduced visibility and heightened noise, which hinder subsequent vision-related tasks [ye2023llod, li2024infrared, xiong2024semantic, lu2025diffdual]. Low-light image enhancement (LLIE) aims to reconstruct perceptually natural scenes by establishing mappings between low-light and normal-light distributions. However, this problem is fundamentally ill-posed, as natural luminance transitions follow continuous physical processes governed by scene radiance and sensor responses, rather than discrete pixel-level correspondences.
While recent deep learning methods—whether supervised [yi2023diff, cai2023retinexformer, lv2018mbllen, shen2017msr, yam2005innovative] or unsupervised [jiang2024lightendiffusion, yang2023difflle, shi2024zero]—attempt to model light variations through paired or unpaired training, they inherently overfit to static relationships between low/normal-light domains. Supervised methods rely on pixel-level correspondences in paired data, forcing models to prioritize localized correlations over the physics of gradual luminance evolution. Unpaired approaches, though avoiding direct pairing, still depend on pseudo-references derived from empirical gamma corrections [jiang2024lightendiffusion], inheriting prior biases. Both paradigms oversimplify the inherently context-dependent and continuous nature of luminance dynamics, resulting in constrained generalization: models excel in domain-specific scenarios with reference normal-light samples but struggle to adapt to unseen environments or sensor-specific degradations [tu2024taming, hou2023global]. This underscores the necessity for a paradigm shift toward learning luminance transitions from the intrinsic continuity of real-world illumination processes.
@@ 2.1 LLIE via Pixel-Level Consistency
Contemporary approaches focusing on pixel-level consistency can be categorized into three evolutionary stages [yi2023diff, wu2023learning, xu2022snr, xu2023low, yang2023implicit, ma2022toward, liu2021retinex, wang2022local]. Early works like LLNet [wei2018deep] and Retinex-Net [zhang2021beyond] leveraged paired datasets to train CNNs with pixel-wise losses, achieving precise local corrections but suffering from domain overfitting. Methods like EnlightenGAN [jiang2021enlightengan] introduced cycle-consistency constraints, while Zero-DCE [guo2020zero] used non-paired training with empirical illumination curves, both inheriting biases from heuristic priors [li2024infrared]. Recent diffusion models [jiang2024lightendiffusion] enhanced flexibility through noise-to-clean transitions, and FeatEnHancer [ye2023llod] proposed hierarchical feature fusion to bridge pixel-level and semantic gaps. These methods improved generalization but remained black-box mappings [xiong2024semantic]. While progressively reducing dependency on strict pixel correspondences, these methods universally prioritize pixel-wise fidelity over the continuous, context-dependent nature of luminance transitions, resulting in constrained generalization when facing unseen scenarios or sensor-specific degradations [tu2024taming].
@@ 2.2 LLIE with Illumination Priors
To alleviate these issues, several studies [guo2020zero, wang2024unsupervised, wang2023low] have incorporated illumination-aware correction into deep neural networks, framing LLIE as a curve-estimation problem through different gamma correction samples. Specifically, [wang2023low] introduced learnable gamma transforms for illumination adjustment, yet enforced uniform corrections across all pixels. Methods like KinD [wang2024unsupervised] decomposed images into illumination-reflectance components, but relied on manually designed reflectance priors that oversimplified real-world radiance interactions, and its reliance on downstream pre-training objectives (e.g., perceptual losses) further introduces external biases that impair adaptability to diverse degradation patterns. More latest works like LightenDiffusion [jiang2024lightendiffusion] integrated Retinex theory into diffusion steps, while GPP-LLIE [zhou2025low] embedded illumination gradients as diffusion guidance. Though enhancing physical plausibility, these methods still imposed global correction strategies through rigid equation constraints.
@@ 3.3 Hierarchically‐Guided Diffusion
Forward Process with Hierarchical Guidance: The sampled set $\mathcal{H}=\{\mathcal{I}_{H}^{(n)}\}_{n=1}^{N}$ employs stochastic learning via diffusion transitions, exploiting the Markov property—each $\mathcal{I}_{H}^{(n)}$ relies only on its forerunner—to adaptively direct noise injection. The low-light image $\mathcal{I}_{L}$ and $\mathcal{H}$ are encoded together using $\mathcal{E}(\cdot)$, incorporating $k$ residual blocks and max-pooling for latent features $\mathcal{F}_{L}\in\mathbb{R}^{\frac{H}{2^{k}}\times\frac{W}{2^{k}}\times C}$ and $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}\in\mathbb{R}^{N\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}\times C}$. We align the $T$-step diffusion with $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}$ using a temporal mapping $\psi:\{1,\ldots,T\}\to\{1,\ldots,N\}$, $N\leq T$, by $\psi(t)=\lfloor t\cdot N/T\rfloor$, such that:
$T=\sum_{n=1}^{N}|T_{n}|,\quad t\in T_{n}\Rightarrow x_{0}=\mathcal{F}_{H}^{(\psi(t))}.$ (7)
The forward diffusion adds Gaussian noise progressively:
$q(x_{t}|x_{t-1})=\mathcal{N}\!\big(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I\big),$ (8)
where $t\in\{1,\dots,T\}$ denotes the diffusion timestep, $x_{t}$ represents the random variable at timestep $t$, and $\beta_{t}$ denotes the noise variance. Accordingly, for each temporal interval $T_{n}$, the corresponding spatial variant $\mathcal{F}_{H}^{(\psi(t))}$ is utilized as the illumination normalization reference, thereby maintaining luminance-consistent forward sampling. By incrementally incorporating spatial luminance awareness from coarse to fine scales into the diffusion forward trajectory, the model acquires a multi-level representation of illumination dynamics. This hierarchical perception facilitates adaptive noise scheduling and enhances robustness across a wide range of lighting conditions.
@@
where the $\mathcal{{D}}(\cdot)$ denotes the decoder. During inference, under the guidance of the low-light input $\mathcal{F}_{L}$, we employ the diffusion model’s implicit sampling strategy [song2020denoising] for reverse denoising. The model utilizes its learned distribution to fit the optimal illumination-enhanced feature representation $\hat{\mathcal{F}}_{N}$, which is then decoded to yield the final output $\hat{\mathcal{I}}_{N}=\mathcal{D}(\hat{\mathcal{F}}_{N})$.
@@ 4.1 Experimental Settings
All experiments are carried out on a cluster of four NVIDIA A800 GPUs under Python 3.9 and PyTorch 2.0, with a fixed batch size of $16$. We employ the Adam optimizer [kingma2014adam], setting the denoising diffusion process learning rate to $2\times 10^{-5}$, while using a sampling ratio $k=3$. The hyperparameters $\lambda_{\text{d}}$, $\lambda_{\text{g}}$ and $\lambda_{\mathrm{GAN}}$ (if activated) are set to $0.9$, $0.005$ and $0.7$ respectively. Noise estimation during diffusion training is performed using the U-Net [ronneberger2015u] architecture with $T=1000$ time steps.
@@ 4.2 Datasets and Metrics
We validated our approach using both paired and unpaired low-light benchmarks. For the paired evaluation, we used the LOLv1 [wei2018deep] and LSRW [hai2023r2rnet] test sets—each comprising matched low- and normal-illumination image pairs—and reported restoration fidelity via PSNR and SSIM [wang2004image], alongside the full-reference perceptual score LPIPS [zhang2018unreasonable]. To assess performance in the absence of ground-truth references, we then tested on the unpaired LIME [guo2016lime], DICM [lee2013contrast], NPE [wang2013naturalness], and VV [vonikakis2018evaluation] collections, measuring perceptual quality with the no-reference NIQE [mittal2012making] and PI [blau20182018] metrics. Our comparative study encompasses six supervised approaches (RetinexNet [wei2018deep], KinD++ [zhang2021beyond], LCDPNet [wang2022local], URetinexNet [wu2022uretinex], SMG [xu2023low] and PyDiff [zhou2023pyramid]) and six unsupervised approaches (Zero-DCE [guo2020zero], EnlightenGAN [jiang2021enlightengan], SCI [ma2022toward], PairLIE [fu2023learning], SCL-LLE [liang2022semantically], LightenDiffusion [jiang2024lightendiffusion] and NeRCo [yang2023implicit].
@@ 4.3 Qualitative Comparison
As shown in Fig. 3, LASQ achieves enhanced local brightness adaptation and superior detail fidelity comparable to supervised methods URetinexNet [wu2022uretinex] and KinD++ [zhang2021beyond] on ground truth-annotated datasets, while demonstrating improved domain-adaptive color reproduction through integration with unpaired normal-light images in LASQ++. By contrast, existing methods exhibit distinct limitations: SCI [ma2022toward] and SCL-LLE [liang2022semantically] suffer from persistent underexposure, whereas EnlightenGAN [jiang2021enlightengan] produces blurred structural details, and NeRCo [yang2023implicit] tends to generate localized over-exposure artifacts. Furthermore, Fig. 4 confirms LASQ’s exceptional performance in real-world scenarios through its complete avoidance of local overexposure, noise amplification, and artifacts that persistently affect other methods: EnlightenGAN [jiang2021enlightengan], NeRCo [yang2023implicit] and PairLIE [fu2023learning] notably show severe localized overexposure and lens flare artifacts, and even supervised approaches URetinexNet [wu2022uretinex] and KinD++ [zhang2021beyond] still struggle to fully suppress these issues. Crucially, LASQ maintains natural scene characteristics without compromising detail fidelity or color consistency, thereby demonstrating unprecedented cross-scenario generalization capability across both constrained laboratory settings and unconstrained environmental conditions. This comprehensive evaluation systematically validates LASQ’s technical superiority in terms of adaptive illumination control, artifact suppression, and domain transfer effectiveness. More results will be provided in the Appendix.
@@ 4.4 Quantitative Comparison
The quantitative evaluation results across diverse datasets are summarized in Table 1, where LASQ demonstrates performance parity with leading supervised techniques on LOLv1 [wei2018deep] and LSRW [hai2023r2rnet] while achieving state-of-the-art results among unsupervised methods through integration of unpaired normal-light images. Notably, on datasets DICM [lee2013contrast], NPE [wang2013naturalness], and VV [vonikakis2018evaluation], LASQ outperforms existing approaches across most perceptual metrics, thereby confirming its intrinsic generalization prowess without domain-specific adaptation. Although normal-light reference integration improves color fidelity, its tendency toward overfitting partially counteracts the model’s inherent generalization capacity, resulting in metric degradation in LASQ++. Crucially, this performance highlights LASQ’s fundamental advantage in balancing domain adaptation with cross-scenario robustness, whereas LASQ++ prioritizes target-domain color accuracy at the expense of slight generalization capability. The extended experimental results, including comprehensive qualitative analyses and quantitative evaluations, are provided in the Appendix.
@@ 5.1 Fixed Luminance Adjustment
We replace our adaptive MCMC-based luminance adaptation with static, hand‐crafted functions (e.g., global gamma correction or fixed tone curves) drawn from prior LLIE methods [huang2012efficient, kallel2018ct, huang2016adaptive, wang2020experiment]. Embedding these into the forward diffusion degrades PSNR, SSIM, and LPIPS (Table 3), showing reduced feature richness and generative fidelity (Fig. 5). In contrast, our physics‐driven, multi‐scale operator—via power‐law stratification and adaptive sampling—better balances global consistency and local detail, yielding stronger cross‐scenario generalization and perceptual quality.
@@ Appendix A The Detailed LASQ Workflow
Hierarchically-Guided Diffusion: The forward process employs a hierarchically-guided diffusion framework to progressively enhance low-light images through semantically-aligned noise injection. By encoding the input $\mathcal{I}_{L}$ and its luminance-enhanced variants $\mathcal{H}$ into latent features $\mathcal{F}_{L}$ and $\left\{\mathcal{F}_{H}^{(n)}\right\}_{n=1}^{N}$, the method establishes temporal correspondence between diffusion steps and multi-scale guidance via a mapping function $\psi(t)$. This alignment dynamically steers the noise trajectory, where each transition step incorporates hierarchical semantic shifts derived from $\mathcal{F}_{H}^{(\psi(t))}$ to modulate the degradation path. Throughout the forward pass, coarse-to-fine illumination characteristics are gradually imprinted into the diffusion process. During reverse denoising, the network $\epsilon_{\theta}$ iteratively removes noise while synergizing hierarchical cues from original low-light features $\mathcal{F}_{L}$. The denoising trajectory is regularized by a dual objective: a noise prediction loss ensures faithful reconstruction aligned with hierarchical semantics, while a global constraint $\mathcal{L}_{\text{g}}$ enforces the structural consistency.
@@ Appendix B The Algorithmic Framework
13: Define $\psi(t)=\lceil tN/T\rceil$, $x_{0}=\mathcal{F}_{H}^{(0)}$
@@
16:  $x_{0}=\mathcal{F}_{H}^{(\psi(t))}$
17:  $x_{t}\sim\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$